# Video Prediction Policy: A Generalist Robot Policy with Predictive Visual Representations

### v1

@@ 5 Experiments
In this section, we conduct extensive experiments on both simulated and real-world robotic tasks to evaluate the performance of the video prediction policy (VPP). The simulated environments include the CALVIN benchmark [39] and MetaWorld benchmark [57], while the real-world tasks encompass Panda arm manipulation and XHand dexterous hand manipulation. Our aim to answer the following questions:
@@
5.1 Simulated Benchmarks Experiments
@@
Baselines. We mainly consider two types of baselines, methods with direct action learning and methods related to future prediction:
  • RT-1 [7]. A direct action learning robot policy that integrates semantic information using Efficient-Net with FiLM-conditioning, followed by token learners for action learning.
  • Diffusion Policy [16]. A direct action learning policy with novel action diffusers.
  • Robo-Flamingo [32]. A direct action learning policy that leverages a pre-trained LLM, incorporating visual information into each layer in a flamingo style [1].
  • Uni-Pi [18]. Begins by learning a video prediction model to generate future sequences and then learns an inverse kinematics model between two frames to determine actions.
  • MDT [47]. Learns a diffusion transformer policy along with an auxiliary mae loss to reconstruct one masked future frame.
  • Susie [5]. Uses a fine-tuned InstructPix2Pix [9] model to generate a goal image and learns a downstream diffusion policy conditioned on the goal image.
  • GR-1 [51]. Learns video and action sequences jointly using an auto-regressive transformer. During policy execution, GR-1 outputs one future frame followed by one action.
Additionally, we include the 3D Diffuser Actor [30] baseline on the Calvin benchmark, as it is the previous state-of-the-art method on this benchmark, although it additionally uses depth image with camera pose unlike other methods.
@@
Video Prediction Policy Training Details. We first train a controllable text-guided video prediction model for the manipulation domain on various datasets as described in Figure 3. Our experiments include 193,690 human manipulation trajectories from the Something-Something-V2 datasets [20] and 179,074 high-quality trajectories from internet robotic manipulation datasets [42, 19, 7, 28, 17, 40]. This stage also includes downstream task datasets, such as the official Calvin ABC dataset and Metaworld dataset, and self-collected datasets on real-world robots. Given the varying scales and quality of different robot datasets, we apply varying sampling probabilities similar to the approach used in [49]. Detailed dataset scales and sample ratios are available in the Appendix 2. The video model training process takes two days on eight NVIDIA A100 GPUs. Subsequent action learning for each robot takes approximately 6-12 hours on four NVIDIA A100 GPUs.
Video Prediction Policy Execution Details. To enhance the control frequency of robots, we assign most of the parameters to the video former part, which has approximately 300M parameters, while the diffusion policy head contains only 20M parameters. The policy execution involves running the video diffusion model and video former for one forward step, and the lightweight diffusion transformer policy denoises the action for 10 steps conditioned on learnable tokens. This design allows us to run the entire video prediction policy process at 7-10 Hz on a local machine equipped with an NVIDIA RTX-4090 GPU. Following the original diffusion policy paper [16], we also output 6$\sim$10 action steps in one VPP forward step, further improving control frequency.
Quantitative Results. The comparisons on the Calvin benchmark are shown in Table 1. Results for Robo-Flamingo, Susie, GR-1, and 3D Diffuser Actors are recorded from their original papers. The MDT result is run on official implementation. The RT-1 result is sourced from [32] and the Uni-Pi result from [5]. We also ran the Diffusion Policy based on the official open-source codebase with CLIP language conditions. Our proposed Video Prediction Policy significantly improved the previous state-of-the-art result from an average task completion length of 3.35 to 4.29 without using any point cloud or depth input. Even with only 10% of the annotated Calvin ABC data used for training, our method still achieved a length of 3.25, which exceeds the results of related methods using full data. Furthermore, the Video Prediction Policy also achieved the best performance in the MetaWorld benchmark with 50 tasks, outperforming the strongest GR-1 baseline by 10.8% in average success rate.
@@
Figure 5: Visualization of the ground-truth video, the complete denoised video, and one-step forward video predictions. Although the textures and details are not precise in the one-step forward videos, they still provide valuable information on physical evolution.
5.2 Analysis of Predictive Visual Representations
Our video prediction policy has achieved significant improvements in simulated experiments with predictive representations. In this part, we conduct various experiments to verify the effectiveness of these predictive representations.
@@ 5.2 Analysis of Predictive Visual Representations
Visualizations of Predictive Representations. Since we use the video prediction model as a vision encoder and perform a single forward pass to obtain predictive representations, we are curious about the quality of these representations. In Figure 5 , we visualize the ground truth future, single-step predictions, and 30-step denoised predictions. Although the single-step prediction does not capture every detail with perfect accuracy, it still conveys valuable information related to robotic manipulation, such as the movement of objects and the robot arm, which effectively supports downstream action learning.
Bridge VideoFusion Tune-A-Video Seer VPP
[…]
Table 3: Quantitative evaluation of prediction quality on bridge datasets. The results of VideoFusion [35], Tune-A-Video [52], Seer [22] are copied from [22].
@@
Video Prediction
Diffusion Model
Video Generation 4.29
@@
Prediction Quality of Manipulation TVP Model. Additionally, we evaluate the quantitative FVD metric [50] on the bridge datasets [19] with complete 30 steps denoising as in [22]. The results are shown in Table 3. Surprisingly, our model easily outperforms the previous TVP model. We attribute this improvement to our use of the pre-trained video foundation model SVD [6], which the earlier TVP model did not leverage, giving us a significant advantage.
Ablation Type Average Length $\uparrow$
VPP 4.29
VPP w/o Internet data 3.97
@@
1.63
Table 5: Ablation study on video pre-training and internet manipulation datasets.
Figure 6: Influences of layer positions and initial noise scales.
Comparisons with Other Visual Representations. To assess our predictive visual representations, we replaced them with alternative visual representations while maintaining other components of the Video Prediction Policy (VPP) unchanged. We considered visual representations pre-trained for different purposes: (1) Stable-VAE [6] pre-trained with VAE image reconstruction loss; (2) VC-1 [38] pre-trained with masked autoencoder loss, tailored for embodied tasks. According to the original study, we finetuned VC-1 on the Calvin datasets using MAE loss to better adapt to the new domain; (3) Voltron [29] pretrained with both MAE reconstruction and language generation tasks. The results, presented in Table 4, indicate that replacing our predictive visual representations leads to a clear decline in performance.
Figure 7: Two real-world hardware platforms and visualizations of sampled tasks. In the Panda arm platform, our generalist policy solves 30+ tasks in 6 skills. In the Xhand dexterous platform, our generalist policy solves 100+ tasks in 13 skills. Challenging tool-use tasks are rendered in blue.
Effectiveness of Video Pre-training and Internet Manipulation Datasets. A significant advantage of the VPP is its ability to leverage the physical knowledge encoded in pre-trained video generation models and Internet manipulation datasets. We conducted experiments to verify the effectiveness of these two components. As shown in Table 5, removing the co-trained Internet manipulation data resulted in a performance decrease from 4.29 to 3.97. Further removing the pre-trained SVD model and training the video prediction model on the Calvin data from scratch led to a substantial performance decline.
Influence of Layer Position and Initial Noise Scales. We are also interested in how different layers of representation and initial white noise scales influence the predictive representations. We experimented with representations from different upsample layers and various initial white noise by altering the total diffusion time-step $t$, following [53]. The results are shown in Figure 6. Our findings suggest that the most effective predictive representations are located in the middle of the upsample blocks rather than the final prediction pixels. Additionally, the quality of representation is not sensitive to initial noise scales.
@@ 5.3 Real World Experiments
We further verified the Video Prediction Policy on two real-world hardware platforms:
  • Franka Panda Robot Arm. On the Franka panda platform, we collected 2k trajectories for over 30+ tasks of 6 categories including picking, placing, pressing, routing, opening, and closing.
  • Xarm with 12-degree Xhand Dexterous Hand. On the dexterous hand platform, we collected 3k trajectories over 100+ tasks of 13 categories, including picking, placing, cup-upright, relocating, stacking, passing, pressing, unplugging, opening, closing, pouring, suction and knocking. Notabley, we also successfully solve the tool-use tasks which are challenging such as picking hammer to hit the nail, grasping the electrical drill and using pipette to transfer the liquid in chemistry experiments.
Figure 8: Evaluations on real-world seen/unseen tasks.
We employ the same text-guided video prediction (TVP) model as in our simulated experiments, trained on both internet datasets and our self-collected real-world data. We train multi-task generalist policies for the Franka Panda and Xhand Dexterous hands, respectively, to solve all tasks in the domain. The hardware platform and visualizations of some selected tasks are shown in Figure 7.
Quantitative Results. Due to the complexity of deploying methods on real-world hardware, we select the strongest baseline models—GR-1, Susie, and the widely-used diffusion policy—as our baselines. We categorize the tasks into “seen” and “unseen” to assess the model’s capabilities. The unseen tasks include new backgrounds and objects that do not appear in the dataset. For evaluation, we perform 200+ rollouts for Panda arm manipulation tasks and 500+ rollouts for dexterous hand manipulation tasks. Due to space constraints, we report only the average success rate in Figure 8. Detailed success rates can be found in Appendix 1, and videos of the roll-out trajectories are available in the supplementary.
@@ 6 Conclusion
We introduce Video Prediction Policy (VPP), a novel approach for learning a generalist robot policy by leveraging predictive visual representations from a video prediction model. Our results show that the representations generated by video prediction models are highly valuable for robot policy learning, yielding consistent improvements across both simulated and real-world tasks. We aim to highlight the potential of video generation models in embodied tasks and underscore the importance of visual representation learning in developing generalist robot policies.
Supplementary Material
For your convenience, a merged video of our rollouts is included in the supplementary zip file.
1 Real-world experiments
1.1 Panda Maniplation
@@ 1.2 Dexterous Manipulation
Table 7: Specific success rate at category level. In seen tasks, We evaluate pick and place tasks 100 times and other tasks 25 times respectively. In unseen tasks, we evaluate pick and place tasks 50 times and other tasks 20 times respectively
Method Tasks completed in a row
1 2 3 4 5 Avg. Len $\uparrow$
VPP(Ours) 0.957 0.912 0.863 0.810 0.750 4.29
VPP(Single-view) 0.909 0.815 0.713 0.620 0.518 3.58
Ablation.1 0.949 0.900 0.839 0.780 0.714 4.18
Ablation.2 0.951 0.904 0.840 0.777 0.718 4.19
Table 8: More ablation studies.
2 Video Prediction Model
2.1 Datasets Sample Ratios
Given the varying quality and scale of these datasets, we have introduced different sample ratios to appropriately balance the influence of different datasets, similar to [49]. Detailed information is shown in Table 9.
2.2 More Visualization of Complete Prediction Results
We present additional visualizations of prediction results from our fine-tuned manipulation TVP model. Predictions on human manipulation datasets are displayed in Figure 10, and those on robotic manipulation datasets are illustrated in Figure 11. All trajectories are sampled from the validation datasets and are predicted using the same manipulation TVP model. Each sample was denoised in 30 steps using classifier-free guidance set at 7.5, as described in [22]. Our TVP model predicts a horizon of 16, and we visualize 8 frames at a skip step of 2 due to space constraints.
2.3 More Visualizations of Predictive Representations
3 More Details for Experiments
3.1 Structure details
We provide the VPP architecture and hyperparameter setting details in four evaluate environments, as shown in Table 10. The transformer block in TVP follows the setting in [6], and the rest of the hyperparameter in Diffusion Transformer follows the work [47].
3.2 More ablation
[…]
Ablation 1 entails the removal of the Temporal-attn module from the Video Former while maintaining all other configurations same as VPP. The results, displayed in Table 8, demonstrate that the Temporal-attn module could enhance the temporal comprehension capabilities of the Video Former.
Ablation 2 introduces a 2-step denoising process in the TVP to derive the predictive visual representation. The outcomes are summarized in Table 8, revealing that the 2-step process did not yield superior performance. We hypothesize this is because a single denoising step suffices to generate an effective representation for trajectory prediction in our configuration. Additionally, the 2-step denoising process nearly doubles the inference time and reduces the control frequency by half. Due to these factors, we opted for a one-step direct encoder in our main experiments.
[…]
3.3 Baseline Implementations
The baseline methods, including RT-1 [7], GR-1 [51], and Diffusion Policy [16], are implemented based on their official repositories. For comparison with Susie [5] in both the Metaworld and real-world manipulation scenarios, we adopt InstructPix2Pix [9] as the future frame predictor and use an image-goal Diffusion Policy [16] to generate the state sequence.

### v2

@@ 5 Experiments
In this section, we conduct extensive experiments on both simulated and real-world robotic tasks to evaluate the performance of the video prediction policy (VPP). We aim to answer the following questions:
@@
5.1 Simulation Setups and Baselines
@@
MetaWorld Benchmark. Metaworld (Yu et al., 2020) features a Sawyer robot performing various manipulation tasks and is widely used to evaluate the precision and dexterity of robotic policies. As shown on the right of Figure 3, it includes 50 tasks with a rich array of operating objects at different levels of difficulty (Radosavovic et al., 2023). We use official Oracle policy to collect 50 trajectories for each task as our training dataset.
@@
VPP Training Details. As outlined in Sec. 4, we use a two-stage training process. In the first stage, we fine-tune a video foundation model into a manipulation-focused TVP model. The videos used in this stage include 193,690 human manipulation trajectories (Goyal et al., 2017) and 179,074 robotic manipulation trajectories (O’Neill et al., 2023), along with downstream task videos, such as the official Calvin ABC videos, the MetaWorld videos, and real-world videos. Given the varying scales and quality of these datasets, we apply different sampling ratios, following the approach in Octo (Team et al., 2024). Detailed dataset scales and sampling ratios can be found in Appendix B. Fine-tuning the video model takes 2-3 days on eight NVIDIA A100 GPUs. In the second stage, we train a generalist policy with Calvin or Metaworld dataset, which requires approximately 6-12 hours on four NVIDIA A100 GPUs.
@@
Figure 4: Visualization of one-step forward visual representations. We can observe that one-step representation already provide valuable information on physical evolution, although the textures and details are not precise.
Policy Roll-out Details. Previous works choose to denoise high-precision videos, a process that is time-consuming and results in low-frequency (Black et al., 2023), or even open-loop control (Du et al., 2024). In contrast, our approach uses the TVP model as an encoder rather than a denoiser, ensuring that each observation is processed through the TVP model only once, which takes less than 160 ms. Then down-stream policy generate action conditioned on the predictive representation. This modification allows us to achieve a significantly higher frequency of 7-10 Hz with consumer-level NVIDIA RTX 4090 GPU. Additionally, we implement action chunking (Chi et al., 2023) with 10 steps to further improve the control frequency.
Comparisons. Generalist robot policy has been widely explored in previous studies. In our experiments, we opted to compare against a representative subset of prior methods that have either achieved state-of-the-art performance or share a similar approach with our methods.
  • RT-1 (Brohan et al., 2022). A direct action learning robot policy that integrates semantic information using Efficient-Net with FiLM-conditioning, followed by token learners for action learning.
  • Diffusion Policy (Chi et al., 2023). A direct action learning policy with novel action diffusers.
  • Robo-Flamingo (Li et al., 2023). A direct action learning policy that leverages a pre-trained LLM, incorporating visual information into each layer in a flamingo style (Alayrac et al., 2022).
  • Uni-Pi (Du et al., 2024). Begins by learning a video prediction model to generate future sequences and then learns an inverse kinematics model between two frames to determine actions.
  • MDT (Reuss et al., 2024). Learns a diffusion transformer policy along with an auxiliary mae loss to reconstruct one masked future frame.
  • Susie (Black et al., 2023). Uses a fine-tuned InstructPix2Pix (Brooks et al., 2023) model to generate a goal image and learns a downstream diffusion policy conditioned on the goal image.
  • GR-1 (Wu et al., 2023a). Learns video and action sequences jointly using an auto-regressive transformer. During policy execution, GR-1 outputs one future frame followed by one action.
  • Robo-Uniview (Liu et al., 2024). Learns a 3d-aware visual encoder with 3d occupation loss to assist policy learning.
  • Vidman (Wen et al., 2024). Pre-trained on the Open X-Embodiment dataset (OXE) video datasets and use a layer-wise self-attention adapter to transform video representation into policy model. However, Vidman did not finetune video model on down-stream tasks which lead to sub-optimal performance.
Quantitative Results. The comparisons on the Calvin benchmark are shown in Table 1. Results for Robo-Flamingo, Susie, GR-1, and 3D Diffuser Actors are recorded from their original papers. The MDT result is run on official implementation. The RT-1 result is sourced from (Li et al., 2023) and the Uni-Pi result from (Black et al., 2023). We also ran the Diffusion Policy based on the official open-source codebase with CLIP language conditions. Our proposed Video Prediction Policy significantly improved the previous state-of-the-art result from an average task completion length of 3.35 to 4.33. Even with only 10% of the annotated Calvin ABC data used for training, our method still achieved a length of 3.25, which exceeds the results of related methods using full data. Furthermore, the Video Prediction Policy also achieved the best performance in the MetaWorld benchmark with 50 tasks, outperforming the strongest GR-1 baseline by 10.8% in average success rate.
@@
Visualizations of Predictive Representations. Since we use the video prediction model as a vision encoder and perform a single forward pass to obtain predictive representations, we are curious about the quality of these representations. In Figure 4 , we visualize the ground truth future, single-step predictions, and 30-step denoised predictions. We can observe that single-step representation already conveys valuable information, such as the movement of objects and the robot arm, which effectively supports downstream action learning.
5.2 Ablation Study
VPP achieves significant improvements in simulated experiments. In this section, we conduct ablation studies to identify the effectiveness of different components of VPP. All ablation study are performed on Calvin ABC-D benchmark and evaluated with average task completion length.
Effectiveness of Predictive Visual Representations. To verify the effectiveness of representation inside VDM, we replace the VDM vision encoder with several other pre-trained vision encoders designed for embodied tasks, while keeping all other components and settings unchanged.
  1. 1.
    Stable-VAE (Blattmann et al., 2023a), pre-trained with a VAE image reconstruction loss. Since the VAE encoder-decoder already performs well in reconstructing images from video datasets, we did not perform further fine-tuning. The input 256×256 images are encoded into 32×32 features with VAE, which are then resampled into 256 tokens via resampler (Jaegle et al., 2021) before passing to the diffusion policy, consistent with VPP.
  2. 2.
    VC-1 (Majumdar et al., 2023), pre-trained with a masked autoencoder loss. The authors note that fine-tuning vc-1 encoder with MAE loss on downstream task datasets can significantly improve performance. For a fair comparison, we first fine-tuned the model on the same video datasets used in VPP. The vc-1 features are resampled into 256 tokens with resampler and pass to policy head.
  3. 3.
    Voltron (Karamcheti et al., 2023), pre-trained with both MAE future reconstruction and language generation tasks. We also fine-tuned the model on our video datasets and resampled the features into 256 tokens.
The results, presented in Table 3, indicate that replacing our predictive visual representations leads to a clear decline in performance.
@@ 5.2 Ablation Study
VDM (ours) Video Generation 4.33
@@
Ablation Type Avg. Length $\uparrow$ Latency $\downarrow$
VPP 4.33 $\sim$140ms
VPP w/o Internet data 3.97 $\sim$140ms
VPP w/o Calvin video 3.31 $\sim$140ms
@@
1.63 $\sim$140ms
VPP w/o Video Former 3.86 $\sim$450ms
VPP w/o Feature Agg. 3.60 $\sim$140ms
Table 4: Ablation study on video pre-training and architecture.
Effectiveness of Video Pre-training and Internet Manipulation Datasets. A significant advantage of the VPP is its ability to leverage the physical knowledge encoded in pre-trained video generation models and Internet manipulation datasets. We conducted experiments to verify the effectiveness of these two components. As shown in Table 4, removing the co-trained Internet manipulation data resulted in a performance decrease from 4.33 to 3.97. Further removing the pre-trained SVD model and training the video prediction model from scratch on the Calvin dataset led to a substantial performance drop. Notably, removing the video pretraining on Calvin alone also caused a significant decline.
Effectiveness of Video Former. The Video Former module plays a pivotal role in extracting predictive representations from the TVP model. To evaluate its effectiveness, we conduct an ablation study by removing the Video Former and directly connecting the TVP features to the diffusion policy. The results, presented in Table 5, are obtained by evaluating the complete VPP model on a single NVIDIA RTX 4090 GPU. The VPP score decreases from 4.33 to 3.86, while the inference time nearly triples. These findings indicate that the absence of the Video Former leads to a substantial degradation in both accuracy and computational efficiency compared to the full model.
Figure 5: Two real-world hardware platforms and visualizations of sampled tasks. We consider a task as “unseen task” if the operated object or the background scene do not appear in training datasets.
Figure 6: Predictions and executions on unseen tasks. Video prediction model generate reasonable futures (red). Real execution trajectories (green) is also closely aligned to the video predicted future (red).
Effectiveness of Feature Aggregation Module. Many previous works (Black et al., 2023; Wu et al., 2023a) directly use the final predicted image to learn policies. However, the image from the final layer often contains many irrelevant details that are not beneficial for the task. In contrast, we adopt a feature aggregation mechanism to leverage multiple layers of features within the up-sampling layers. We replace aggregated features with final layer features while keeping the other layers unchanged. This process lead to a decrease in the average task completion length on the Calvin benchmark, from 4.33 to 3.60. More ablations on different layers can be found at Appendix C.2.
@@ 5.3 Real World Experiments
We further verified the Video Prediction Policy on two real-world hardware platforms.
Franka Panda Robot Arm. On the Franka Panda platform, we collected 2,000 trajectories for over 30 tasks in 6 categories: picking, placing, pressing, routing, opening, and closing. We divided the tasks into seen and unseen categories. A task is considered unseen if the operated object is new or the background scene is new.
Xarm with 12-degree Xhand Dexterous Hand. On the dexterous hand platform, we collected 4,000 trajectories over 100+ tasks in 13 categories, including picking, placing, cup-upright, relocating, stacking, passing, pressing, unplugging, opening, closing, pouring, suction, and knocking. We also define a task as unseen if the operated object is new or the background scene is new. Additionally, we included four challenging tool-use tasks, including the use of a spoon, hammer, electrical drill, and pipette for chemistry tasks. More task details can be found in Appendix A.
Franka Panda DP Susie GR-1 VPP(ours)
Seen Tasks 0.42 0.56 0.52 0.85
Unseen Tasks 0.25 0.46 0.38 0.73
Dexterous Hand DP Susie GR-1 VPP(ours)
Seen Tasks 0.28 0.45 0.32 0.75
Unseen Tasks 0.11 0.28 0.15 0.60
Tool-use Tasks 0.05 0.23 0.15 0.68
Table 5: Success rates on real-world tasks. Due to space limit, we only show the average success rate on each category. Detailed success rate can be found at Appendix A
Training and Rollout Details. We employ the same text-guided video prediction (TVP) model as in our simulated experiments, trained on both internet datasets and collected real-world data. Then a generalist robot policy is learned to solve all tasks in the domain conditioned on instructions. The hardware platform and visualizations of some selected tasks are shown in Figure 5.
Quantitative Results. Due to the complexity of deploying methods on real-world hardware, we select the strongest baseline models—GR-1, Susie, and the widely-used diffusion policy—as our baselines. For evaluation, we perform 200+ rollouts for Panda arm manipulation tasks and 500+ rollouts for dexterous hand manipulation tasks. The comparisons are in the Table 5, which indicate VPP outperforms all the baselines with a clear margin in both seen tasks, unseen tasks and tool-use tasks.
Generalization Analysis. we take three unseen tasks as case studies: picking up a tennis ball, pouring Coca-Cola, and using a spoon. Notably, none of these objects—tennis ball, Coca-Cola, or spoon—appear in our collected dataset. As illustrated in Figure 6, the video prediction model forecast reasonable future states even on unseen tasks. Moreover, we observe that the actual execution trajectory closely aligns with the predicted future state. We interpret the generalization mechanism of the VPP model in two key aspects: First, video models can make correct visual predictions even on unseen tasks due to internet-scale pre-training; Second, the low-level policy learns a robust inverse dynamics model that only needs to implicitly track the movement of the robot in the predicted future, without the need to focus on new objects or backgrounds. In this way, the VPP model successfully generalizes to a wide range of unseen tasks.
@@ 6 Conclusion
We introduce Video Prediction Policy (VPP), a novel approach for learning a generalist robot policy. VPP learns an implicit inverse dynamics model conditioned on predictive representations inside VDMs and yields consistent improvements across both simulated and real-world tasks. As video generation models are more and more powerful these days, we aim to fully unlock the power of video model in building physical intelligence and highlight the potential of video generation models in embodied tasks.
Appendix A Real-world experiments
A.1 Panda Maniplation
@@ A.2 Dexterous Manipulation
Tool-use Tasks
Diffusion
Policy
Susie GR-1 VPP
Spoon 0.0 0.4 0.3 0.9
Hammer 0.2 0.2 0.1 0.6
Drill 0.0 0.1 0.2 0.8
Pipette 0.0 0.0 0.0 0.4
Average 0.05 0.23 0.15 0.68
Table 7: Specific success rate at category level. In seen tasks, We evaluate pick and place tasks 100 times and other tasks 25 times respectively. In unseen tasks, we evaluate pick and place tasks 50 times and other tasks 20 times respectively. We evaluate each tool-use task for 10 times.
Appendix B Video Prediction Model
B.1 Datasets Sample Ratios
Given the varying quality and scale of these datasets, we have introduced different sample ratios to appropriately balance the influence of different datasets, similar to (Team et al., 2024). Detailed information is shown in Table 8.